OpenStudy (anonymous):

Multiply. √(7)(√(5)-√(3))

OpenStudy (anonymous):

I got √(35-√21 am I correct?

OpenStudy (anonymous):

=sqrt35-sqrt21 = 1.33

OpenStudy (anonymous):

according to the given you are right

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt {7} \left( \sqrt {5}-\sqrt {3} \right) \] \[\Large \sqrt {7}\times\sqrt {5}-\sqrt {7}\times\sqrt {3} \] \[\Large \sqrt {7\times5}-\sqrt {7\times3} \] \[\Large \sqrt {35}-\sqrt {21} \] So \[\Large \sqrt {7} \left( \sqrt {5}-\sqrt {3} \right)= \sqrt {35}-\sqrt {21}\] So you have the correct answer.